The light shining on a diffraction grating has a wavelength of 495 nm (in vacuum). The grating produces a second-order bright fringe whose position is defined by an angle of 9.34. How many lines per centimeter does the grating have?
Answer #1

The condition principle maxima for diffraction by grating is,

d\sin \theta =m\lambda

Here, d is the spacing between the grating lines, \theta is the angle of diffraction, m is the order of the spectrum, and

\lambda is the wavelength of light used.

Rearrange the equation for d.

     d =\frac{m\lambda}{\sin \theta}

         =\frac{2\left (495\times 10^{-9} \textup{ m} \right )}{\sin\left ( 9.34^{\textup{o}} \right )}

        =6.1\times 10^{-6} \textup{ m}

The number of lines per on the grating is,

N=\frac{1}{d}

=\frac{1}{6.1\times 10^{-6} \textup{ m}}

=\frac{1}{\left (6.1\times 10^{-6} \textup{ m} \right )\left ( \frac{1 \textup{ cm}}{0.01 \textup{ m}} \right )}

=1639.32 /\textup{ cm}

Therefore, the number of lines per centimeter on the grating approximately is1640.